OpenStudy (anonymous):

The base of a triangle one-third of the triangle's height. Find the area of the triangle as a function of the base.

OpenStudy (johnweldon1993):

Alright well we know...the area of a triangle is \[\large A = \frac{1}{2}bh\] If we know that the base...is 1/3 of the hieght...we know \[\large b = \frac{1}{3}h\] So if the 'b' = that...we can rewrite the first equation we have up there as \[\large A = \frac{1}{2}(\frac{1}{3}h)h\] Notice all we did was write in \(\large \frac{1}{3}h\) instead of the 'b' we had before

OpenStudy (anonymous):

So what would i have to do next?

OpenStudy (anonymous):

its multiple choice.

OpenStudy (johnweldon1993):

Lets see...what are the choices?

OpenStudy (anonymous):

1. A=3/2 b^2 2. A=9/2 b^2 3. A=1/6 b^2 4. A=1/18 b^2

OpenStudy (johnweldon1993):

Oh they want to replace the hieght instead...alright...so lets start over lol

OpenStudy (johnweldon1993):

Area of triangle \[\large A = \frac{1}{2}bh\] The base is 1/3 the height....also meaning that the height is 3 times the base....so \[\large h = 3b\] If THAT is true...we can now replace 'h' in our original equation...with '3b' \[\large A = \frac{1}{2}b(3b)\] Now we simplify... \[\large A = \frac{1}{2}(3b^2)\] \[\large A = \frac{3}{2}b^2\]

OpenStudy (anonymous):

thank you soooo much :*

OpenStudy (johnweldon1993):

Anytime ;)